OpenStudy (anonymous):

Please healp solve: 18x^2-29x=0

OpenStudy (anonymous):

divide by x.... just kidding you'll lose an anser if you do that factor out an x what do you get

OpenStudy (anonymous):

18x-29=0

OpenStudy (anonymous):

x=0 and x=29/18

OpenStudy (anonymous):

nah you can't divide by x because you'll lose an answer

OpenStudy (anonymous):

you'll lose the answer x=0

OpenStudy (anonymous):

and you don't want that so factor it out x(18x-29)=0 x=0 18x-29=0 18x=29 x=29/18 the reason why you don't divide by x is because you're ruling out that x cannot = 0 because you cannot divide by 0

OpenStudy (anonymous):

I understand now.. x=0 and x=29/18

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

just a tip for incase you didn't understand why not divide by it

OpenStudy (anonymous):

if it said x>0 then you could